# Learning More Effective Representations for Dense Retrieval through Deliberate Thinking Before Search

### v1

@@ Abstract
Recent dense retrievers usually thrive on the emergency capabilities of Large Language Models (LLMs), using them to encode queries and documents into an embedding space for retrieval. These LLM-based dense retrievers have shown promising performance across various retrieval scenarios. However, relying on a single embedding to represent documents proves less effective in capturing different perspectives of documents for matching. In this paper, we propose Deliberate Thinking based Dense Retriever (Debater), which enhances these LLM-based retrievers by enabling them to learn more effective document representations through a step-by-step thinking process. Debater introduces the Chain-of-Deliberation mechanism to iteratively optimize document representations using a continuous chain of thought. To consolidate information from various thinking steps, Debater also incorporates the Self Distillation mechanism, which identifies the most informative thinking steps and integrates them into a unified text embedding. Experimental results show that Debater significantly outperforms existing methods across several retrieval benchmarks, demonstrating superior accuracy and robustness. All codes are available at https://github.com/OpenBMB/DEBATER.
Learning More Effective Representations for Dense Retrieval through Deliberate Thinking Before Search
Yifan Ji1thanks:   indicates equal contribution., Zhipeng Xu [^1] footnotemark: 1, Zhenghao Liu1thanks:   indicates corresponding author., Yukun Yan2, Shi Yu2, Yishan Li3, Zhiyuan Liu2, Yu Gu1, Ge Yu1, Maosong Sun2 1Department of Computer Science and Technology, Northeastern University, China 2Department of Computer Science and Technology, Institute for AI, Tsinghua University, China Beijing National Research Center for Information Science and Technology, China 3ModelBest.Inc
1 Introduction
@@
Dense retrieval models encode both queries and documents into a dense embedding space and measure their similarity to retrieve relevant documents Karpukhin et al. (2020); Zhao et al. (2024); Xiong et al. (2021), demonstrating strong effectiveness in various downstream NLP tasks, such as open-domain question answering Chen and Yih (2020), fact verification Liu et al. (2020), and web search Chen et al. (2024). However, recent findings have shown that dense retrievers suffer from significant performance degradation when applied to new tasks or domains Su et al. (2023), raising concerns about their versatility Luo et al. (2024); Khramtsova et al. (2024).
@@
Figure 1: The Illustration of Our Deliberate Thinking based Dense Retriever (Debater). Debater leverages the reasoning capability of LLM to conduct fine-grained document representations for retrieval.
@@
Large Language Models (LLMs), such as ChatGPT Achiam et al. (2023) and LLaMA Touvron et al. (2023), have demonstrated extraordinary emergent capabilities Wei et al. (2022a); Zhao et al. (2023), inspiring researchers to leverage them to enhance the task and domain generalization of dense retrievers Zhu et al. (2023); Khramtsova et al. (2024). In particular, existing work has focused on prompting LLMs to generate dense representations for retrieval Zhuang et al. (2024). These methods typically use task-specific instructions or in-context demonstrations to guide LLMs in generating task- and domain-aware embeddings. To learn more tailored representations for dense retrieval, researchers further focus on optimizing LLM-based retrievers using relevance labels Ma et al. (2024); Neelakantan et al. (2022); Li et al. (2025). These methods exploit the superior reasoning abilities of LLMs, achieving impressive performance across various retrieval tasks Wang et al. (2024a); Zhu et al. (2023); Luo et al. (2024). Recent studies suggest that LLMs pose strong reasoning capability, particularly implemented by their step-by-step thinking Kudo et al. (2024); Wei et al. (2022b). LLM-based retrievers typically rely on the hidden state of the end-of-sequence token as both query and document representations. Nevertheless, only relying on one embedding usually shows less effectiveness in representing documents from different views that can match queries Zhang et al. (2022); Khattab and Zaharia (2020).
@@
In this paper, we propose a Deliberate Thinking based Dense Retriever (Debater) model to learn more effective document representations through deliberately thinking step-by-step before retrieval. As shown in Figure 1, our method stimulates LLMs to conduct the reasoning process, enabling them to generate more fine-grained document representations for retrieval. Specifically, Debater introduces the Chain-of-Deliberation mechanism to encourage LLMs to conduct deliberate thinking by autograssively decoding the document representations. Then Debater utilizes the Self Distillation mechanisms to gather all information from previous steps and compress them into the document embedding at the last step.
@@
Our experiments show that Debater achieves comparable or even better retrieval performance than the baseline methods implemented by larger-scale LLMs, highlighting its effectiveness. Our further analyses show that both Chain-of-Deliberation and Self Distillation play important roles in Debater and appropriately increasing the thinking steps can benefit these LLM-based dense retrieval models. Thriving on autograssively decoding different document representations during thinking, the document representations can be gradually refined to be more effective. By incorporating our Self Distillation, LLMs show the ability to capture different key information at different thinking steps and gather all crucial semantics from different steps to the final document representations.
2 Related Work
@@
Dense retrieval Karpukhin et al. (2020); Xiong et al. (2021); Su et al. (2023) has proven effective in various NLP downstream tasks Liu et al. (2020); Chen et al. (2024); Guu et al. (2020). However, the versatility of dense retrievers remains a challenge that hinders their progress Luo et al. (2024); Lee et al. (2024), particularly their inability to generate task- and domain-specific embeddings and return suitable results Su et al. (2023); Luo et al. (2024); Tao et al. (2024). To address this limitation, prior work has focused on conducting fine-grained data curation to fine-tune dense retrievers with multi-task instructions Su et al. (2023); Asai et al. (2023). However, obtaining high-quality relevance labels can be difficult for training dense retrievers Yu et al. (2022); Gao et al. (2023); Wang et al. (2024a).
@@
Recent research has shifted towards using LLMs as the backbone for dense retrievers Tao et al. (2024), thriving on their strong emergence capabilities. Some studies attempt to directly prompt LLMs to generate embeddings for retrieval Zhuang et al. (2024). However, prompt-based approaches cannot leverage pre-existing retrieval signals, limiting their effectiveness Zhu et al. (2023). In contrast, recent efforts have focused on fine-tuning LLMs for dense retrieval tasks Wang et al. (2024a); Ma et al. (2024); Li et al. (2024), or designing additional pretraining tasks to transform LLMs into dense retrievers BehnamGhader et al. (2024), achieving strong retrieval performance and generalization capabilities. However, existing methods typically extract the last hidden state of the end-of-sequence token as the dense representation Ma et al. (2024); Luo et al. (2024), which is not always effective for fully representing documents from different perspectives to match queries Zhang et al. (2022); Khattab and Zaharia (2020). The exploration of different document representations, such as leveraging the reasoning ability of LLMs, remains an underexplored area.
@@
Figure 2: The Architecture of Deliberate Thinking based Dense Retriever (Debater). Debater uses both the Chain-of-Deliberation (CoD) and Self Distillation (SD) for training LLMs.
@@
To enhance the reasoning capability of LLMs, one approach is to generate intermediate reasoning steps using Chain-of-Thought (CoT) Wei et al. (2022b) or its variants Chen et al. (2023); Zhang et al. (2024). CoT allows LLMs to delay final answers by engaging in reasoning Kudo et al. (2024), improving response accuracy Wei et al. (2022b); Chu et al. (2024). However, these approaches operate within the language space and often require generating tens or even hundreds of additional tokens, which can hinder their ability to meet the latency requirements of dense retrievers. Current research is exploring the integration of CoT reasoning into a continuous latent space Hao et al. (2024); Xie et al. (2024) to enhance computational efficiency. Building on these advancements, our Debater focuses on latent reasoning chains, encouraging LLM-based retrievers to think step-by-step to enhance the dense representations of documents.
3 Methodology
@@
3.1 Preliminary of Dense Retrieval with Large Language Models as Foundations
@@
$f(q,d)=sim(h^{q},h^{d}),$ (2)
where sim denotes the similarity function. In Debater, we use cosine similarity to measure the similarity between queries and documents, which is also employed in previous works Wang et al. (2024a); BehnamGhader et al. (2024). Subsequently, we train the LLMs contrastively to maximize the probability of retrieving the positive document $d^{+}$ over the negative document $d^{-}$:
$p(d^{+}|q,{d^{+}}\cup\mathcal{D}^{-})=\frac{e^{f(q,d^{+})}}{e^{f(q,d^{+})}+\sum_{d^{-}\in\mathcal{D}^{-}}e^{f(q,d^{-})}},$ (3)
where $\mathcal{D}^{-}$ denotes the set of negative documents, typically obtained via in-batch negative sampling Karpukhin et al. (2020).
@@
3.2 Enhancing Dense Retriever through Deliberate Thinking
@@
$\mathcal{L}_{c}=-\text{log}\frac{e^{f_{\text{max}}(q,{d^{+}})}}{e^{f_{\text{max}}(q,{d^{+}})}+\sum_{d^{-}\in\mathcal{D}^{-}}e^{f_{\text{max}}(q,{d^{-}})}}.$ (7)
@@
$P(d_{i}|q)=\frac{e^{f_{\text{max}}(q,d_{i})}}{\sum_{d_{j}\in\tilde{\mathcal{D}}}e^{f_{\text{max}}(q,d_{j})}},$ (8)
@@
$P(\tilde{\mathcal{D}}|q)=\left[P(d_{1}|q),P(d_{2}|q),\dots,P(d_{k}|q)\right].$ (9)
Each value $P(d_{i}|q)$ represents the ranking probability of the $i$-th document $d_{i}$ using the document representations from all thinking steps $\{h^{d}_{1},...,h^{d}_{m}\}$ of CoD that yield a higher similarity with the query. Concurrently, we compute the rank probability of $d_{i}$ using the last-token embedding $h^{d}_{m}$ from CoD:
$Q(d_{i}(t_{m})|q)=\frac{e^{f(q,d_{i}(t_{m}))}}{\sum_{d_{j}\in\{d^{+}\}\cup\mathcal{D}^{-}}e^{f(q,d_{j}(t_{m}))}}.$ (10)
@@
$Q(\tilde{\mathcal{D}}|q)=\left[Q(d_{1}|q),Q(d_{2}|q),\dots,Q(d_{k}|q)\right].$ (11)
@@
$\mathcal{L}_{t}={P(\tilde{\mathcal{D}}|q)\cdot\log\frac{P(\tilde{\mathcal{D}}|q)}{Q(\tilde{\mathcal{D}}|q)}},$ (12)
@@
$\mathcal{L}=\mathcal{L}_{c}+\mathcal{L}_{t},$ (13)
@@
4 Experimental Methodology
In this section, we describe the datasets, baselines, and implementation details for our experiments.

### v2

# Learning Refined Document Representations for Dense Retrieval via Deliberate Thinking

@@ Abstract.
Recent dense retrievers increasingly leverage the robust text understanding capabilities of Large Language Models (LLMs), encoding queries and documents into a shared embedding space for effective retrieval. However, most existing methods represent each document with a single embedding, which is less effective at capturing its multifaceted semantics and thereby limits matching accuracy. In this paper, we propose Deliberate Thinking based Dense Retriever (Debater), a novel approach that enhances document representations by incorporating a step-by-step thinking process. Debater introduces a Chain-of-Deliberation mechanism, which iteratively refines document embeddings through a continuous chain-of-thought. To integrate information from various thinking steps, Debater further employs a Self Distillation mechanism that identifies and fuses the most informative steps into a unified embedding. Experimental results show that Debater significantly outperforms existing methods across several retrieval benchmarks, demonstrating superior accuracy and robustness. All codes and datasets are available at https://github.com/OpenBMB/DEBATER.
Dense Retrieval, Large Language Models, Deliberate Thinking, Knowledge Distillation
journalyear: 2025ccs: Information systems Information retrieval
1. Introduction
@@
Dense retrieval models encode both queries and documents into a dense embedding space and measure their similarity to retrieve relevant documents (Karpukhin et al., 2020; Zhao et al., 2024; Xiong et al., 2021), demonstrating strong effectiveness in various downstream NLP tasks, such as open-domain question answering (Chen and Yih, 2020), fact verification (Liu et al., 2020), and web search (Chen et al., 2024). However, recent findings have shown that dense retrievers suffer from significant performance degradation when applied to new tasks or domains (Su et al., 2023), raising concerns about their versatility (Luo et al., 2024; Khramtsova et al., 2024).
@@
Figure 1. The Illustration of Our Deliberate Thinking based Dense Retriever (Debater). Debater leverages the reasoning capability of LLM to conduct fine-grained document representations for retrieval.
@@
Large Language Models, such as ChatGPT (Achiam et al., 2023) and LLaMA (Touvron et al., 2023), have demonstrated extraordinary emergent capabilities (Wei et al., 2022a; Zhao et al., 2023), inspiring researchers to leverage them to enhance the task and domain generalization of dense retrievers (Zhu et al., 2023; Khramtsova et al., 2024). In particular, existing work has focused on prompting LLMs to generate dense representations for retrieval (Zhuang et al., 2024). These methods typically use task-specific instructions or in-context demonstrations to guide LLMs in generating task- and domain-aware embeddings. To learn more tailored representations for dense retrieval, researchers further focus on optimizing LLM-based retrievers using relevance labels (Ma et al., 2024; Neelakantan et al., 2022; Li et al., 2025). These methods exploit the superior reasoning abilities of LLMs, achieving impressive performance across various retrieval tasks (Wang et al., 2024a; Zhu et al., 2023; Luo et al., 2024). Recent studies suggest that LLMs pose strong reasoning capability, particularly implemented by their step-by-step thinking (Kudo et al., 2024; Wei et al., 2022b). LLM-based retrievers typically rely on the hidden state of the end-of-sequence token as both query and document representations. Nevertheless, only relying on one embedding usually shows less effectiveness in representing documents from different views that can match queries (Zhang et al., 2022; Khattab and Zaharia, 2020).
@@
In this paper, we propose a Deliberate Thinking based Dense Retriever (Debater) model to learn more effective document representations through deliberately thinking step-by-step before retrieval. As shown in Figure 1, our method stimulates LLMs to conduct the reasoning process, enabling them to generate more fine-grained document representations for retrieval. Specifically, Debater introduces the Chain-of-Deliberation mechanism to encourage LLMs to conduct deliberate thinking by autoregressively decoding the document representations. Then Debater utilizes the Self Distillation mechanisms to gather all information from previous steps and compress them into the document embedding at the last step.
@@
Our experiments demonstrate that Debater achieves retrieval performance comparable to, or even surpassing, that of baseline methods implemented by larger-scale LLMs, demonstrating its effectiveness. Further analysis reveals that Chain-of-Deliberation and Self-Distillation play complementary roles in Debater, and that increasing the number of reasoning steps appropriately can benefit LLM-based dense retrieval models. The document representations produced by Debater are progressively refined through iterative thinking, where each step autoregressively generates intermediate representations. By incorporating Self-Distillation, the model is able to extract different salient information at different reasoning steps and integrate them into a more comprehensive and semantically rich final representation.
2. Related Work
@@
Dense retrieval (Karpukhin et al., 2020; Xiong et al., 2021; Su et al., 2023) has proven effective in various NLP downstream tasks (Liu et al., 2020; Chen et al., 2024; Guu et al., 2020). However, the versatility of dense retrievers remains a challenge that hinders their progress (Luo et al., 2024; Lee et al., 2024), particularly their inability to generate task- and domain-specific embeddings and return suitable results (Su et al., 2023; Luo et al., 2024; Tao et al., 2024). To address this limitation, prior work has focused on conducting fine-grained data curation to fine-tune dense retrievers with multi-task instructions (Su et al., 2023; Asai et al., 2023). However, obtaining high-quality relevance labels can be difficult for training dense retrievers (Yu et al., 2022; Gao et al., 2023; Wang et al., 2024a).
@@
Recent research has shifted towards using LLMs as the backbone for dense retrievers (Tao et al., 2024), thriving on their strong emergence capabilities. Some studies attempt to directly prompt LLMs to generate embeddings for retrieval (Zhuang et al., 2024). However, prompt-based approaches cannot leverage pre-existing retrieval signals, limiting their effectiveness (Zhu et al., 2023). In contrast, recent efforts have focused on fine-tuning LLMs for dense retrieval tasks (Wang et al., 2024a; Ma et al., 2024; Li et al., 2024), or designing additional pretraining tasks to transform LLMs into dense retrievers (BehnamGhader et al., 2024), achieving strong retrieval performance and generalization capabilities. However, existing methods typically extract the last hidden state of the end-of-sequence token as the dense representation (Ma et al., 2024; Luo et al., 2024), which is not always effective for fully representing documents from different perspectives to match queries (Zhang et al., 2022; Khattab and Zaharia, 2020). The exploration of different document representations, such as leveraging the reasoning ability of LLMs, remains an underexplored area.
@@
Figure 2. The Architecture of Deliberate Thinking based Dense Retriever (Debater).
@@
To enhance the reasoning capability of LLMs, one approach is to generate intermediate reasoning steps using Chain-of-Thought (CoT) (Wei et al., 2022b) or its variants (Chen et al., 2023; Zhang et al., 2024). CoT allows LLMs to delay final answers by engaging in reasoning (Kudo et al., 2024), improving response accuracy (Wei et al., 2022b; Chu et al., 2024). However, these approaches operate within the language space and often require generating tens or even hundreds of additional tokens, which can hinder their ability to meet the latency requirements of dense retrievers. Current research is exploring the integration of CoT reasoning into a continuous latent space (Hao et al., 2024; Xie et al., 2024) to enhance computational efficiency. Building on these advancements, our Debater focuses on latent reasoning chains, encouraging LLM-based retrievers to think step-by-step to enhance the dense representations of documents.
3. Methodology
@@
3.1. Preliminary of Dense Retrieval with Large Language Models as Foundations
@@
(2) $f(q,d)=sim(h^{q},h^{d}),$
where sim denotes the similarity function. In Debater, we use cosine similarity to measure the similarity between queries and documents, which is also employed in previous works (Wang et al., 2024a; BehnamGhader et al., 2024). Subsequently, we contrastively train the LLM to maximize the probability of retrieving the positive document $d^{+}$ over the negative document $d^{-}$:
(3) $p(d^{+}|q,{d^{+}}\cup\mathcal{D}^{-})=\frac{e^{f(q,d^{+})}}{e^{f(q,d^{+})}+\sum_{d^{-}\in\mathcal{D}^{-}}e^{f(q,d^{-})}},$
where $\mathcal{D}^{-}$ denotes the set of negative documents, typically obtained via in-batch negative sampling (Karpukhin et al., 2020).
@@
3.2. Enhancing Dense Retriever through Deliberate Thinking
@@
(7) $\mathcal{L}_{c}=-\text{log}\frac{e^{f_{\text{max}}(q,{d^{+}})}}{e^{f_{\text{max}}(q,{d^{+}})}+\sum_{d^{-}\in\mathcal{D}^{-}}e^{f_{\text{max}}(q,{d^{-}})}}.$
@@
(8) $P(d_{i}|q)=\frac{e^{f_{\text{max}}(q,d_{i})}}{\sum_{d_{j}\in\tilde{\mathcal{D}}}e^{f_{\text{max}}(q,d_{j})}},$
@@
(9) $P(\tilde{\mathcal{D}}|q)=\left[P(d_{1}|q),P(d_{2}|q),\dots,P(d_{k}|q)\right].$
Each value $P(d_{i}|q)$ represents the ranking probability of the $i$-th document $d_{i}$ using the document representations from all thinking steps $\{h^{d}_{1},...,h^{d}_{m}\}$ of CoD that yield the highest similarity with the query. Concurrently, we compute the rank probability of $d_{i}$ using the last-token embedding $h^{d}_{m}$ from CoD:
(10) $Q(d_{i}(t_{m})|q)=\frac{e^{f(q,d_{i}(t_{m}))}}{\sum_{d_{j}\in\{d^{+}\}\cup\mathcal{D}^{-}}e^{f(q,d_{j}(t_{m}))}}.$
@@
(11) $Q(\tilde{\mathcal{D}}|q)=\left[Q(d_{1}|q),Q(d_{2}|q),\dots,Q(d_{k}|q)\right].$
@@
(12) $\mathcal{L}_{t}={P(\tilde{\mathcal{D}}|q)\cdot\log\frac{P(\tilde{\mathcal{D}}|q)}{Q(\tilde{\mathcal{D}}|q)}},$
@@
(13) $\mathcal{L}=\mathcal{L}_{c}+\mathcal{L}_{t},$
@@
4. Experimental Methodology
In this section, we describe the datasets, evaluation metrics, baselines, and implementation details for our experiments.
Datasets. We train all Debater models using the public portion of the E5 dataset (Wang et al., 2024a; Springer et al., 2025), a carefully curated collection of approximately 1.5M publicly available samples. Table 1 presents its statistics, and the full dataset is accessible on their website [^1] https://github.com/jakespringer/echo-embeddings.
Table 1. Data Statistics of E5 Dataset. We show the composition and distribution of E5 Dataset.
Dataset #Samples Proportion
ELI5 (Fan et al., 2019) 32,547 2.16%
HotpotQA (Yang et al., 2018) 90,447 5.99%
FEVER (Thorne et al., 2018) 101,578 6.73%
MIRACL (Zhang et al., 2023) 32,561 2.16%
MSMARCO Passage Ranking (Bajaj et al., 2016) 249,592 16.53%
MSMARCO Document Ranking (Bajaj et al., 2016) 73,400 4.86%
NQ (Kwiatkowski et al., 2019) 100,231 6.64%
NLI (Gao et al., 2021) 277,230 18.36%
SQuAD (Rajpurkar et al., 2016) 87,599 5.80%
TriviaQA (Joshi et al., 2017) 73,346 4.86%
Quora Duplicate Questions (DataCanary et al., 2017) 101,762 6.74%
Mr-TyDi (Zhang et al., 2021) 48,715 3.23%
DuReader (He et al., 2018) 86,395 5.72%
T2Ranking (Xie et al., 2023) 154,294 10.22%
Table 2. Data Statistics of the BEIR Benchmark. We show the task type, along with the number of queries and passages for each dataset.
Dataset Task #Query #Corpus
TREC-COVID Bio-Medical 50 171,332
NFCorpus Information 323 3,633
NQ Question 3,452 2,681,468
HotpotQA Answering 7,405 5,233,329
FiQA-2018 (QA) 648 57,638
ArguAna Argument 1,406 8,674
Touché-2020 Retrieval 49 382,545
CQADupStack Duplicate-Question 13,145 457,199
Quora Retrieval 10,000 522,931
DBPedia Entity-Retrieval 400 4,635,922
SCIDOCS Citation-Prediction 1,000 25,657
FEVER Fact Checking 6,666 5,416,568
Climate-FEVER Wikipedia 1,535 5,416,593
SciFact Scientific 300 5,183
The retrieval effectiveness of Debater is evaluated on the BEIR benchmark (Thakur et al., 2021), which includes 18 datasets that span a variety of domains. Our evaluation focuses on the 14 publicly available datasets used for the zero-shot retrieval task. The statistics for these datasets are provided in Table 2.
Evaluation Metrics. To evaluate the retrieval effectiveness of Debater, we use nDCG@10, the standard metric for the BEIR benchmark. The metric implementation follows the pytrec-eval toolkit (Gysel and de Rijke, 2018), which is consistent with prior work (Zhu et al., 2023).
Baselines. We compare Debater with several baseline retrievers implemented with different language models. GTR (Ni et al., 2022) employs large dual encoder-only models to build a dense retriever, while SGPT (Muennighoff, 2022) trains dense retrieval models using decoder-only architectures. Emb-V3222https://cohere.com/blog/introducing-embed-v3 is a commercial text retrieval model provided by Cohere. PromptReps (Zhuang et al., 2024) directly prompts LLMs to generate dense representations without supervision. RepLLaMA (Ma et al., 2024) and E5-Mistral (Wang et al., 2024b) fine-tune LLMs as dense retrievers, using the hidden state of an additional end-of-sequence token to represent the input context. Notably, E5-Mistral is trained on the same dataset as Debater but leverages a larger foundational model.
Table 3. Overall Retrieval Performances on BEIR Benchmark. indicates the $1$ most representative BEIR tasks used in CPT (Neelakantan et al., 2022) and Avg CPT Sub reflects the average performance across these tasks.